OpenStudy (gabylovesyou):

To rent a certain meeting room, a college charges a reservation fee of $36 and an additional fee of $6.60 per hour. The chemistry club wants to spend less than $88.80 on renting the meeting room. What are the possible amounts of time for which they could rent the meeting room? Use t for the number of hours the meeting room is rented, and solve your inequality for t .

OpenStudy (gabylovesyou):

@mathstudent55

OpenStudy (mathstudent55):

room rental = fixed cost + hourly cost ok so far?

OpenStudy (gabylovesyou):

yes

OpenStudy (mathstudent55):

The fixed cost is the reservation fee, so we can replace "fixed cost" with 36 to get room rental = 36 + hourly fee still ok?

OpenStudy (gabylovesyou):

ok

OpenStudy (mathstudent55):

The hourly fee is $6.60/hour If an hour costs $6.60, what do t hours cost?

OpenStudy (gabylovesyou):

t*6.60

OpenStudy (mathstudent55):

Correct. We can write it simply as 6.6t

OpenStudy (gabylovesyou):

ok

OpenStudy (mathstudent55):

Now we have the following cost expression: total cost = 36 + 6.6t Ok?

OpenStudy (gabylovesyou):

ok

OpenStudy (mathstudent55):

They want to spend less than $88. We set our cost expression less than 88 to get our inequality. 36 + 6.6t < 88

OpenStudy (mathstudent55):

How do we know we use < and not \(\le\)? The problem tells us they want to spend LESS than $88. That means the cost must be less than $88, and < means less than. If the problem stated they want to spend AT MOST $88, then they could spend $88 or less, and $88 would have to be included, then you'd use \(\le\). Ok?

OpenStudy (gabylovesyou):

ok

OpenStudy (mathstudent55):

The next step is to solve the inequality.

OpenStudy (gabylovesyou):

ok....36 + 6.6t < 88 t <7.87878787

OpenStudy (mathstudent55):

36 + 6.6t < 88 6.6t < 52 t < 7.8787... You are correct. If they need to rent it for a whole number of hours, then they can rent it for less than 8 hours.